# Influencer Identification on Link Predicted Graphs

### v2

@@ V.1 Results on a Random Graph
V.2 Results on an Empirical Dataset
The application of our algorithm to the arXiv collaboration network led to the following conclusions:
  • our similarity metric RA22-2- 2 has the lowest average MSE overall,
  • the Social Sphere Model may be able to identify latent influencers who perform even better than those in the original graph, and
  • there always exist parameters for our modified centrality algorithms that perform better than the originals.
Prediction Metric Complex Simple Overall
Common Neighbors 0.02641 0.00557 0.01599
Jaccard 0.02282 0.00493 0.01388
Local Path 0.04713 0.00811 0.02762
Quasi-Local RA 0.04289 0.00753 0.02521
Quasi-Local RA-2 0.04081 0.00738 0.02410
RA-2 0.02254 0.00635 0.01444
Resource Allocation 0.02278 0.00543 0.01411
Overall 0.03220 0.00647 0.01933
Table 7: Average MSE over all top-$k$ algorithms for link prediction metrics using $70\%$ training graphs from the arXiv GrQc dataset.
Prediction Metric Complex Simple Overall
Common Neighbors 0.01914 0.01368 0.01641
Jaccard 0.01854 0.01334 0.01594
Local Path 0.02607 0.01365 0.01986
Quasi-Local RA 0.02539 0.01392 0.01966
Quasi-Local RA-2 0.02558 0.01371 0.01965
RA-2 0.01804 0.01372 0.01588
Resource Allocation 0.01846 0.01414 0.01630
Overall 0.02160 0.01374 0.01767
Table 8: Average MSE over all top-$k$ algorithms for link prediction metrics using $90\%$ training graphs from the arXiv GrQc dataset.
Centrality Metrics Common Neighbors Jaccard Local Path Quasi-local RA-2 Quasi-local RA RA-2 Resource Allocation Overall
Balanced Index
0.0151
0.0106
0.0292
0.0260
0.0247
0.0104
0.0116
0.0182
Betweenness
0.0110
0.0094
0.0224
0.0205
0.0194
0.0087
0.0092
0.0144
Closeness
0.0170
0.0138
0.0295
0.0262
0.0249
0.0133
0.0135
0.0197
ClusterRank
0.0174
0.0151
0.0301
0.0296
0.0287
0.0146
0.0149
0.0215
Complex Path Centrality
0.0475
0.0480
0.0442
0.0434
0.0431
0.0486
0.0484
0.0462
Degree
0.0107
0.0097
0.0229
0.0219
0.0195
0.0086
0.0084
0.0145
Eigenvector
0.0109
0.0095
0.0240
0.0223
0.0212
0.0091
0.0098
0.0152
H-index
0.0156
0.0152
0.0318
0.0282
0.0268
0.0150
0.0152
0.0211
k-core
0.0117
0.0103
0.0222
0.0189
0.0187
0.0109
0.0107
0.0148
LeaderRank
0.0108
0.0083
0.0229
0.0211
0.0196
0.0076
0.0080
0.0140
LocalRank
0.0094
0.0080
0.0231
0.0216
0.0205
0.0077
0.0079
0.0140
PageRank
0.0130
0.0087
0.0266
0.0240
0.0237
0.0099
0.0094
0.0165
Overall
0.0158
0.0139
0.0274
0.0253
0.0242
0.0137
0.0139
0.0192
Table 9: Average MSE for centrality metrics over all algorithms and contagion models on $70\%$ training graphs.
Centrality Metrics Common Neighbors Jaccard Local Path Quasi-local RA-2 Quasi-local RA RA-2 Resource Allocation Overall
Balanced Index
0.0132
0.0132
0.0221
0.0221
0.0218
0.0129
0.0132
0.0169
Betweenness
0.0114
0.0107
0.0199
0.0192
0.0198
0.0102
0.0111
0.0146
Closeness
0.0123
0.0117
0.0201
0.0198
0.0197
0.0111
0.0116
0.0152
ClusterRank
0.0120
0.0120
0.0206
0.0202
0.0204
0.0120
0.0118
0.0156
Complex Path Centrality
0.0732
0.0732
0.0376
0.0374
0.0377
0.0733
0.0734
0.0580
Degree
0.0107
0.0102
0.0172
0.0169
0.0170
0.0099
0.0101
0.0131
Eigenvector
0.0077
0.0074
0.0120
0.0117
0.0116
0.0069
0.0072
0.0092
H-index
0.0125
0.0126
0.0182
0.0179
0.0177
0.0122
0.0124
0.0148
k-core
0.0030
0.0029
0.0048
0.0045
0.0046
0.0031
0.0029
0.0037
LeaderRank
0.0115
0.0108
0.0179
0.0175
0.0181
0.0104
0.0105
0.0138
LocalRank
0.0107
0.0109
0.0198
0.0193
0.0195
0.0103
0.0105
0.0144
PageRank
0.0122
0.0116
0.0204
0.0199
0.0199
0.0113
0.0117
0.0153
Overall
0.0159
0.0156
0.0192
0.0189
0.0190
0.0153
0.0155
0.0170
Table 10: Average MSE for centrality metrics over all algorithms and contagion models on $90\%$ training graphs.
Centrality Metric
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Balanced Index
0.3217
0.4166
0.3091
0.3714
0.3691
0.4086
0.4046
0.3716
Betweenness
0.3811
0.3903
0.3549
0.3537
0.3503
0.3937
0.3943
0.3740
Closeness
0.2897
0.3371
0.2720
0.3120
0.3154
0.3463
0.3411
0.3162
ClusterRank
0.2794
0.2731
0.2217
0.1891
0.1806
0.2760
0.2714
0.2416
Complex Path Centrality
0.2171
0.2160
0.1326
0.1320
0.1343
0.2171
0.2166
0.1808
Degree
0.3634
0.3520
0.3400
0.3211
0.3211
0.3594
0.3594
0.3452
Eigenvector
0.2691
0.2154
0.2577
0.2091
0.2057
0.2240
0.2189
0.2286
H-index
0.1846
0.1509
0.0617
0.0571
0.0560
0.1509
0.1497
0.1158
k-core
0.4680
0.4623
0.4417
0.4394
0.4371
0.4634
0.4611
0.4533
LeaderRank
0.3514
0.3491
0.3297
0.3177
0.3154
0.3560
0.3531
0.3389
LocalRank
0.3000
0.2886
0.2200
0.2057
0.2017
0.2977
0.3000
0.2591
PageRank
0.3766
0.3651
0.3503
0.3291
0.3263
0.3674
0.3674
0.3546
Overall
0.3169
0.3180
0.2743
0.2698
0.2678
0.3217
0.3198
0.2983
Table 11: Average overall accuracies for centrality metrics on $70\%$ training graphs from the arXiv GrQc dataset.
Centrality Metric
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Balanced Index
0.4166
0.4869
0.3989
0.4589
0.4571
0.4857
0.4851
0.4556
Betweenness
0.4743
0.4703
0.4366
0.4343
0.4349
0.4709
0.4697
0.4558
Closeness
0.3874
0.3983
0.3714
0.3857
0.3846
0.4006
0.4000
0.3897
ClusterRank
0.2937
0.2743
0.2080
0.1960
0.1903
0.2743
0.2771
0.2448
Complex Path Centrality
0.1680
0.1640
0.1560
0.1566
0.1571
0.1720
0.1680
0.1631
Degree
0.4606
0.4206
0.4229
0.3869
0.3869
0.4223
0.4194
0.4171
Eigenvector
0.3789
0.3131
0.3600
0.2977
0.3029
0.3217
0.3189
0.3276
H-index
0.1834
0.1743
0.0863
0.0869
0.0834
0.1714
0.1726
0.1369
k-core
0.5274
0.5217
0.4909
0.4829
0.4857
0.5211
0.5211
0.5073
LeaderRank
0.4640
0.4326
0.4274
0.4006
0.4006
0.4360
0.4349
0.4280
LocalRank
0.3280
0.3194
0.2246
0.2200
0.2183
0.3194
0.3194
0.2784
PageRank
0.4851
0.4480
0.4497
0.4143
0.4177
0.4509
0.4486
0.4449
Overall
0.3806
0.3686
0.3360
0.3267
0.3266
0.3705
0.3696
0.3541
Table 12: Average overall accuracies for centrality metrics on $90\%$ training graphs from the arXiv GrQc dataset.
Mean Squared Error Comparison. Considering Table 9 and Table 10, one can see that the mean squared error of the percentage infected is still small, though substantially larger than before, likely due to the larger variation that comes with more vertices and a larger graph. Notably, our new link prediction metric has the smallest average MSE over all the prediction metrics in both scenarios, with it also being the smallest in $75\%$ of the metrics in Table 10; we also see a similar trend in Table 11 and Table 12, where our new metric has the highest fraction in the $70\%$ case and the second highest in the $90\%$ scenario. These results support our hypotheses, as there is a significant decrease in MSE in the $90\%$ scenario as opposed to the $70\%$ one, this indicating that our link prediction metrics can identify influencers with very similar influence to that of the original, thus effectively showing prediction.
Accuracy Comparison. When one compares the two scenarios described in Table 11 and Table 12, one can see more variation in the influencers chosen in the second scenario due to less information; in the $90\%$ case, the graph is already very similar to the final configuration and link prediction is more accurate, whereas predictions may be less grounded for the latter case. For more specific data, see Appendix A.
(a) On complex contagion
(b) On simple contagion
Figure 6: Average performance of each algorithm over all graphs as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset.
(a) On complex contagion
(b) On simple contagion
Figure 7: Average performance of each algorithm over all graphs as fraction infected over time using $90\%$ training graphs from the arXiv GrQc dataset.
Influence Comparison. In what follows, we shall consider the relative influences of influencers selected by the following methods:
  1. 1.
    the algorithms,
  2. 2.
    the various prediction metrics when controlling for algorithm, and
  3. 3.
    the various centrality metrics when controlling for algorithm.
We have mapped out the average performances of the various algorithms in Figure 6 and Figure 7. Notice that, as in the random graph case, VoteRank performs admirably in both simple and complex contagion in each scenario. Intriguingly, the $k$ highest algorithm also seems to do well, particularly in complex contagion. Unlike in Section V.1, LIR and LIR22-2- 2 exhibit sub-optimal efficiency and have effectively swapped with graph coloring and joint nomination. A possible explanation for this observation is the difference in structure of the empirical dataset and the Erdös-Rényi random graph. For instance, the empirical dataset is not fully connected (and it is thus impossible for the fraction infected to be $1$, as there are too many disconnected pieces).
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 8: Average performance of each algorithm and prediction metric as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset in complex contagion.
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 9: Average performance of each algorithm and prediction metric as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset in simple contagion.
When one controls for algorithm and investigate the average influences of the various prediction metrics (see Figure 8 and Figure 9), we can see that
  • in most cases the respective influence plots of the original graph and the various predicted graphs are very close, and
  • influencers found on the predicted graphs are often able to spread their influence to more of the population.
The first observation entails that the Social Sphere Model is able to produce results comparable with influence identification on the original graph. To elaborate on the second, scenarios such as in 7(a), 7(d), and 8(e) depict the influence plots of predicted graphs as eventually plateauing to values above those of the original graph (and the training graph). This suggests that link predicted graphs may sometimes be able to find latent influencers, those that were influential in the past but overlooked in the current graph.
One should note that Xie et al. dynamicedgedecay have previously explained the drawbacks to having a model that only considers a small window of the graph’s historical data: a powerful influencer who is perhaps temporarily inactive may be overlooked in favor of a currently active, less influential individual. Importantly, our Social Sphere Model thus also presents an incorporation of historical data in finding influencers.
Finally, our analysis of the effectiveness of our modifications to the various algorithms showed that for VoteRank, there exist centrality VoteRank metrics whose averages in accuracy or MSE over all the predicted graphs match or surpass VoteRank’s performance in both the $70\%$ and $90\%$ scenarios, namely closeness, complex path, H-index, $k$-core, and LocalRank in the $70\%$ graph and Balanced Index and LeaderRank in the $90\%$ graph.
When comparing the relative influence of the nodes chosen with VoteRank over the various link prediction metrics, the above is also true (see Figure 10, Figure 11, and Appendix A); in particular, betweenness is shown to be one of the best performing metrics in each scenario (None refers to VoteRank) and closeness and complex path centrality also often do almost as well. This is likely because our contagion models are deterministic and are concerned with the distances between vertices, which the three mentioned metrics, being path-based, are geared to do. Similar observations hold for LIR and LIR22-2- 2.
(a) Original Graph
(b) Common Neighbors Predicted Graph
(c) deg2 RA Predicted Graph
(d) Jaccard Predicted Graph
(e) Local Path Predicted Graph
(f) Quasi-Local deg2 ra Predicted Graph
(g) Quasi-Local RA Predicted Graph
(h) Resource Allocation Predicted Graph
Figure 10: Performance of influencers chosen with Centrality VoteRank (the original VoteRank is labelled as None) on predicted graphs formed from $70\%$ training graphs from the arXiv GrQc dataset in complex contagion.
(a) Original Graph
(b) Common Neighbors Predicted Graph
(c) Jaccard Predicted Graph
(d) Local Path Predicted Graph
(e) Quasi-Local RA Predicted Graph
(f) Quasi-Local RA-2 Predicted Graph
(g) RA-2 Predicted Graph
(h) Resource Allocation Predicted Graph
Figure 11: Performance of influencers chosen with Centrality VoteRank (the original VoteRank is labelled as None) on predicted graphs formed from $70\%$ training graphs from the arXiv GrQc dataset in simple contagion.
VI Conclusion
In this research, we introduced the Social Sphere Model, a novel algorithm which utilizes heuristic link prediction and influencer identification to forecast future influencers in weighted networks. Our comprehensive analyses on both a random graph and a real world scenario (through the arXiv collaboration network) have demonstrated the algorithm’s efficacy, particularly evident in the low mean squared error and relatively high accuracy, suggesting the Social Sphere Model’s promising prediction capabilities for detecting future influencers, which is similar to Yanchenko et al. exante ’s findings that the influence of their model’s selected nodes were comparable to that of the Oracle method.
@@ VI Conclusion
  • Viral marketing: in the scenario mentioned in Section I, taking our algorithm over existing social media data and estimating $t$ from the expected time for production can assist in identifying potential future influencers in the network, giving more time (a valuable commodity) for negotiation, scouting, and strategizing. The relatively short time complexities of heuristic models are also sometimes valuable in these scenarios. As an analogy, if one wanted to understand how admissions would look like in a university program for influencers, our algorithm would allow one to evaluate the potential of an influencer or group of influencers in the wider network.
  • Better Network Approximation: The already-documented crimeprediction utility of link prediction in approximating incomplete network data further underscores the significance of our approach, especially considering the typically resource-intensive nature of complete data collection.
@@
In particular, our model differentiates from other ones by making use of relatively low time complexity metrics and algorithms as opposed to neural networks, the training of which may be infeasible for smaller businesses that might seek to calculate future influencers. The simplicity and efficiency of it can thus allow for quick identification of influencers and thus better networking.
Since our model is inspired by a deterministic simple contagion model, which restricts the differentiation in algorithm performance, we present here an interesting future research path to explore more realistic contagion models and extend the Social Sphere Model further.
Finally, it would be of much interest to study the implications of the Social Sphere Model more graph theoretically on different network types: in particular, towards identifying the most effective combinations of link prediction and influencer identification metrics for specific network types, such as temporal or directed networks, varying empirical dataset origins, and properties like varying clustering coefficient or diameter values.
Acknowledgments. We extend our gratitude to the MIT PRIMES-USA program for their invaluable support and guidance throughout this research. The research of LPS is partially supported by NSF FRG Award DMS- 2152107, a Simons Fellowship and NSF CAREER Award DMS 1749013. Part of the work appearing here was done whilst LPS was a Visiting Fellow at All Souls College, Oxford.
Affiliations.
  1. (a)
    University of Illinois at Chicago, Chicago, IL 60607, USA.
  2. (b)
    Lakeside High School, Seattle, WA 98107.
Appendix A More Data
Figure 14: Average accuracies for different algorithms on $70\%$ training graphs from the arXiv GrQc dataset.
Figure 15: Average accuracies for different algorithms on $90\%$ training graphs from the arXiv GrQc dataset.
A complete set of simulation data, csv files, and line plots can be found at https://drive.google.com/drive/folders/1WyuwbMddzajawt4nfh7rLrztY99zO6sm?usp=sharing.

### v3

@@ 5 Results
6 Conclusion
In this research, we introduced the Social Sphere Model, a novel algorithm which utilizes heuristic link prediction and influencer identification to forecast future influencers in weighted networks. Our comprehensive analyses on a complex random graph have demonstrated the algorithm’s efficacy, particularly evident in the low mean squared error and relatively high accuracy, suggesting the Social Sphere Model’s promising prediction capabilities for detecting future influencers, which is similar to Yanchenko et al. [65]’s findings that the influence of their model’s selected nodes were comparable to that of the Oracle method.
@@ 6 Conclusion
  • Viral marketing: in the scenario mentioned in Section 1, taking our algorithm over existing social media data and estimating $t$ from the expected time for production can assist in identifying potential future influencers in the network, giving more time (a valuable commodity) for negotiation, scouting, and strategizing. The relatively short time complexities of heuristic models are also sometimes valuable in these scenarios. As an analogy, if one wanted to understand how admissions would look like in a university program for influencers, our algorithm would allow one to evaluate the potential of an influencer or group of influencers in the wider network.
  • Better Network Approximation: The already-documented [6] utility of link prediction in approximating incomplete network data further underscores the significance of our approach, especially considering the typically resource-intensive nature of complete data collection.
@@
In particular, our model differentiates from other ones by making use of relatively low time complexity metrics and algorithms as opposed to neural networks, the training of which may be infeasible for smaller businesses that might seek to calculate future influencers. The simplicity and efficiency of it can thus allow for quick identification of influencers and thus better networking. Finally, it would be of much interest to study the implications of the Social Sphere Model more empirically on various synthetic networks and real-world datasets such as Yelp, Wikipedia, Coauthorship Network, and Twitter while investigating properties like varying clustering coefficient or diameter values.
Finally, whilst this paper focuses on the theoretical formulation of the Social Sphere Model, including its integration of similarity-based link prediction, and evaluation of the model on random graphs, a companion work will examine the empirical performance and practical applications of the model on empirical datasets. Our second study, titled "Empirical Evaluation of the Social Sphere Model on Influencer Identification on Social Networks," we evaluate the model’s behavior in a practice setting on author collaboration networks, and its predictive ability in identifying latent influencers, and explore the application for information diffusion and digital marketing strategies. The two works put together aim to provide a comprehensive investigation of the Social Sphere Model’s capabilities from a theoretical lens and in real-world datasets.
Acknowledgments
We extend our gratitude to the MIT PRIMES-USA program for their invaluable support and guidance throughout this research. The research of LPS is partially supported by NSF FRG Award DMS- 2152107, a Simons Fellowship and NSF CAREER Award DMS 1749013. Part of the work appearing here was done whilst LPS was a Visiting Fellow at All Souls College, Oxford.
[…]